OpenStudy (anonymous):

2/x-1 = 3/x+2 . The answer is is 7 however I do not understand how they got this answer >:(

OpenStudy (anonymous):

Multiply throughout by x to get 2-x=3+2x => 3x=-1 => x=-1/3. The answer is not 7.

OpenStudy (anonymous):

The answer could be 7. 2/(x-1) = 3/(x+2), x = 7. (2/x)-1 = (3/x)+2, x = -1/3.

OpenStudy (anonymous):

for 2/(x-1) = 3/(x+2): cross multiply: 2(x+2) = 3(x-1) expand terms: 2x + 4 = 3x -3 add -3x -4 to both sides: -x = -7 divide by -1: x = 7

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hahaha. Quantish. I did not notice that.